OpenStudy (anonymous):

How do I find anything in problem out because I feel like there are missing components....the problem is: A car covers 72 kilometers in the first hour of its journey. In the next hour, it covers 90 kilometers. What is the amount of the work done by the car? The total mass of the car, including its passengers, is 2.5 * 10^3 kilograms. This is really confusing.

OpenStudy (anonymous):

Sorry to ask such a long and multi-stepped problem.....but I am just soo confused by it....and thx for taking your time and answering....:D

OpenStudy (anonymous):

First of all ,if the car wasnt going up a hill or anything, the work done by gravity is 0. Then maybe, since the car has an average velocity of (72+90)/2 =81 kmh^-1, it has a kinetic energy and the gain in kinetic is work done by the car's engine.Not so sure though

OpenStudy (anonymous):

so the mass doesn't have anything to do with the problem?

OpenStudy (anonymous):

ya it does ... the convert 81 km per hour into meter per second ,and kinetic energy = 1/2*m*v^2. m=total mass and v* velocity. 81 km per hour is 22.5 meter per second

OpenStudy (anonymous):

like i have five options..... A. 2.8 * 10^5 B. 4.0 * 10^5 C. 5.1 * 10^5 D. 2.8 * 10^6 E. 2.8 * 10^7

OpenStudy (anonymous):

if u did what i said u will get the answer as 6.32*10^5 ... so i might be wrong

OpenStudy (anonymous):

ok well thx for the help honestly u helped a lot no sarcasm at all so thx